my problem is: Luke has $5.00 more than Sam, together they have $73.00. How much does each have. Explain the equation.
I know the answer is $34.00 and $39.00, but i don't know how to write it out.
Thank you

Found 3 solutions by stanbon, checkley71, Earlsdon:
Answer by stanbon(48568)

Luke has $5.00 more than Sam, together they have $73.00. How much does each have. Explain the equation.
--------------
Start word problems by symbolizing the unknows.
Let amount Sam has be "x" dollars.
Then Luke has "x+5" dollars.
---------------------------
Then try to write an equation involving the variables.
Sam + Luke have 73 dollars
x + (x+5) = 73
Then solve for the variable.
2x+5=73
2x= 68
x=34
Then interpret the results.
Sam has x=34 dollars
Luke has x+5=34+5=39 dollars
--------------
Hope this helps.
Cheers,
Stan H.

Answer by checkley71(8403)

X+(X+5)=73
X+X+5=73
2X=73-5
2X=68
X=68/2
X=34 ANSWER FOR SAM
34+5=39 FOR LUKE
PROOF
34+39=73
73=73

Answer by Earlsdon(6103)

Try this!
Let L = the amount that Luke has and S = the amount that Sam has. From the problem description, you can write:
L = S+$5 and
L+S = $73 Substitute here the L from above and solve for S.
(S+$5) + S = $73 Simplify.
2S+$5 = $73 Subtract $5 from both sides.
2S = $68 Divide both sides by 2.
S = $34
L = S+$5
L = $39